OpenStudy (vishweshshrimali5):

I just thought of some problem: Find out the value of: \(1!*2!*3!*...*n!\)

OpenStudy (kirbykirby):

I believe this is called the "superfactorial"

OpenStudy (vishweshshrimali5):

I tried simplifying it: \(1!*2!*3!*...*n!\) \[\prod_{i = 1}^{n} (n+1-i)!\] \[\prod_{i=1}^n (i)^{n+1-i}\] But how to solve it after this ? Any ideas ?

OpenStudy (vishweshshrimali5):

sorry no idea about that @kirbykirby , it was just some random question that came in my mind.

OpenStudy (kirbykirby):

I don't think it simplifies any more

OpenStudy (vishweshshrimali5):

Well I think that's the end of another idea :(

OpenStudy (vishweshshrimali5):

Thanks @kirbykirby

OpenStudy (vishweshshrimali5):

How about this one: \[\large{\prod_{k=1}^{n}(k!)^{\sum_{i=1}^{k}(i!)}}\]

OpenStudy (kirbykirby):

There may be some weird way of writing it in another way though But I'm not aware of it. I only came across the superfactorial once so I don't know any nice properties for it

OpenStudy (vishweshshrimali5):

Any idea ?

OpenStudy (kirbykirby):

would it not simply be \(\prod_{k=1}^n k!\) ?

OpenStudy (vishweshshrimali5):

No I don't think so.

OpenStudy (vishweshshrimali5):

Lets change it into this: \(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 How about this one: \[\large{\prod_{k=1}^{n}(k!)^{\large{\sum_{i=1}^{k}(i*i!)}}}\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (vishweshshrimali5):

This one is much easier I think: \(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 Lets change it into this: \(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 How about this one: \[\large\color{red}{\prod_{k=1}^{n}(k!)^\color{blue}{\large{\sum_{i=1}^{k}(i*i!)}}}\] \[= \large\color{red}{\prod_{k=1}^{n}(k!)^\color{blue}{\large{(k+1)!-1}}}\]

OpenStudy (vishweshshrimali5):

Now any idea how to continue this ?

OpenStudy (vishweshshrimali5):

What do you think @kirbykirby

OpenStudy (vishweshshrimali5):

I used this property: \[\large{\sum_{i=1}^{k} (i!)} = (k+1)! - 1 \]

OpenStudy (vishweshshrimali5):

@mathmale

OpenStudy (kirbykirby):

I don't see why \[\prod_{k=1}^n k!\] is not appropriate. Maybe k! should be in parentheses to be clear. \[ \prod_{k=1}^n (k!)\] .

OpenStudy (vishweshshrimali5):

There was no problem in that. I was thinking that you were telling me the net solution for the problem. Sorry about that. :)

OpenStudy (kirbykirby):

Oh you were giving different problem..

OpenStudy (kirbykirby):

sorry I didn't realize

OpenStudy (vishweshshrimali5):

I am getting this on further simplification: \[\large{\prod_{k=1}^{n} (k)^{\sum_{i=1}^{n}[(i+1)!-1]}}\]

OpenStudy (vishweshshrimali5):

Lets solve the power first. That will make the problem less messy.

OpenStudy (vishweshshrimali5):

Okay, I have, \[\large{\sum_{i=1}^{n}[(i+1)!-1]}\]

OpenStudy (vishweshshrimali5):

I believe I cracked it

OpenStudy (kirbykirby):

I'm not aware of that identity, and it doesn't seem to work \[\large{\sum_{i=1}^{k} (i!)} = (k+1)! - 1 \] say k = 4: LHS 1! + 2! + 3! + 4! = 33 RHS: 5! - 1 = 119

OpenStudy (vishweshshrimali5):

Oh sorry I meant this: \[\large{\sum_{i=1}^{k} (i*i!)} = (k+1)! - 1 \]

OpenStudy (kirbykirby):

oh ok that makes sense

OpenStudy (vishweshshrimali5):

Sorry about that

OpenStudy (kirbykirby):

Are you looking for a closed-form solution

OpenStudy (vishweshshrimali5):

Well I am frankly looking for the most simplified form

OpenStudy (vishweshshrimali5):

Solving the power: Okay, I have, \[\large{\sum_{i=1}^{n}[(i+1)!-1]}\] \[=\large{\sum_{i=1}^{n}[(i+1)!]-n}\] Again stuck here. How to find the sum of factorials?

OpenStudy (kirbykirby):

I don't think this is a simple solution: http://mathworld.wolfram.com/FactorialSums.html

OpenStudy (vishweshshrimali5):

Yeah but that was if n was not a natural number

OpenStudy (vishweshshrimali5):

That's why gamma function was used.

OpenStudy (vishweshshrimali5):

http://math.stackexchange.com/questions/451604/can-the-factorial-function-be-written-as-a-sum This one is much better

OpenStudy (vishweshshrimali5):

But I don't think its going to raise our spirits even a small bit.

OpenStudy (vishweshshrimali5):

I think the best thing we can say is just leave it. Its really frustrating.

OpenStudy (vishweshshrimali5):

Another one of my idea into the bin. :D

OpenStudy (kirbykirby):

but if you are doing a summation from 1 to n in sigma notation, then it should be natural numbers. Otherwise you would be doing some integral if you could allow any number over an interval

OpenStudy (kirbykirby):

but I agree dealing with factorials is not obvious

OpenStudy (vishweshshrimali5):

Yeah... I think I am going to change my ideas to calculus. At least, it doesn't have such things. On second thought, I am not so sure about it too. But, I will surely give it a try. I remember one thought: Always try before giving up. So that you can say that at least I got defeated after trying.

OpenStudy (vishweshshrimali5):

Well thanks for sticking with me @kirbykirby . I really appreciate your help. Good day.

OpenStudy (kirbykirby):

well good effort on your part, maybe you will discover a nice formula sometime :)

OpenStudy (vishweshshrimali5):

May be...